# InspectorRAGet: An Introspection Platform for RAG Evaluation

### v1

@@ Abstract
Large Language Models (LLM) have become a popular approach for implementing Retrieval-Augmented Generation (RAG) systems, and a significant amount of effort has been spent on building good models and metrics. In spite of increased recognition of the need for rigorous evaluation of RAG systems, few tools exist that go beyond the creation of model output and automatic calculation. We present InspectorRAGet, an introspection platform for RAG evaluation. InspectorRAGet allows the user to analyze aggregate and instance-level performance of RAG systems, using both human and algorithmic metrics as well as annotator quality. InspectorRAGet is suitable for multiple use cases and is available publicly to the community1. The demo video is available at https://youtu.be/MJhe8QIXcEc
@@ 1 Introduction
The development of all RAG systems have one thing in common: the need for rigorous evaluation. Evaluating such systems requires continuously designing experiments including datasets, models, and metrics; running evaluations; and analyzing the results to address shortcomings and make appropriate deployment decisions. Businesses wishing to deploy a RAG system must evaluate it on their specific use cases (not only academic benchmarks), often comparing multiple architectures and models.
Recognizing the importance of evaluation, the research community has been creating evaluation benchmark datasets Liu et al. (2023); Chen et al. (2023), evaluation metrics Es et al. (2023), and noise robustness Chen et al. (2023), as well as evaluation frameworks (such as RAGAs Es et al. (2023), and ARES Saad-Falcon et al. (2023)). Leveraging these assets, practitioners can automatically compute aggregate evaluation scores for their RAG systems. However, this provides only a very limited lens of a system’s performance. We suggest that a comprehensive evaluation of RAG systems should include the following:
Aggregate Performance: Overall evaluation is important for continuous benchmarking of models and dataset performance Gehrmann et al. (2022).
Instance-level Analysis: Aggregate metrics alone do not offer much insight into the RAG system, particularly in identifying the source of undesirable output. A flexible, feature-rich workflow for detecting and inspecting related individual instances empowers the researcher to perform actionable error analysis.
Mixed Metrics: Algorithmic metrics such as BLEU and ROUGE do not necessarily align with human judgements Reiter (2018); Blagec et al. (2021); nor are automatic evaluations using LLMs Es et al. (2023) a perfect substitute. However, a comprehensive analysis of all types of metrics can yield a rich array of insights.
Annotator Qualification: Even human judgements of RAG systems are imperfect Chiang and Lee (2023). A thorough understanding of annotator behavior allows the researcher to identify and improve ambiguous guidelines, complex data points, and underperforming annotators, resulting in higher quality evaluations.
Dataset Characterization: The dataset itself should be subjected to a thorough inspection during evaluation. Fixing erroneous reference answers, clarifying ambiguous instances, or even identifying bias in the content can improve the overall evaluation outcome by providing much needed context for the observed quantitative results.
Achieving such a multi-pronged evaluation would currently require researchers to manually examine their evaluation output, perhaps creating various ad-hoc data processing scripts and visualizations in spreadsheet tools. Our goal is to significantly lower the required effort for this important and undersupported aspect of RAG evaluation. Users should be able to easily drill down from aggregate performance overviews to instance-level views of data, models, metrics and annotators, in order to perform detailed error analysis and identify actionable insights.
Our contributions are as follows:
@@
We present InspectorRAGet, an interactive evaluation platform with a rich feature set that addresses each of the desired aspects of a comprehensive RAG evaluation: performance benchmarking, a combined aggregate and instance level analysis, a holistic view of results via a mix of metrics, annotator qualification, and dataset characterization.
@@
We use InspectorRAGet to yield concrete and actionable insights on two use cases, on both new and existing datasets.
@@
We open source the platform for use by the community111https://github.com/IBM/InspectorRAGet.
@@ 2 The RAG Evaluation Life Cycle
We begin by briefly describing the life cycle of evaluating RAG systems, and situating InspectorRAGet within it, as shown in Figure 1.
Evaluating RAG systems involves three main steps:
(1) Design the evaluation experiment: One starts by designing the evaluation experiment (or simply experiment), which consists of Models, Datasets, Metrics and Metric Evaluators, as defined below.
(2) Run the evaluation experiment: Once the experiment has been designed, the next step is to compute the evaluation scores. This involves (a) computing the model responses and (b) passing the responses to the metric evaluators (algorithms or humans) to produce instance-level scores.
(3) Analyze the experiment results: Finally, the experiment results need to be analyzed to gain actionable insights about the models, datasets, metrics, annotation quality, etc.
@@ 2.1 InspectorRAGet for RAG Evaluation
InspectorRAGet focuses on the third step of the RAG evaluation life cycle. To analyze an experiment using the platform, model developers or stakeholders upload a standardized JSON file summarizing the experiment results. This summary contains the following information:
@@
• Metric metadata: Metadata about the metrics on which the model responses were evaluated. These include the metric name, type (i.e., algorithmic, if it was evaluated by an algorithm or an LLM used as a judge and human, if it was evaluated by humans, such as crowd workers), and scale (yes/no, Likert scale, numeric, etc.).
• Evaluation scores: The evaluation scores of models along each metric on every data instance 222In case of multiple annotators per metric, such as when multiple humans annotate a single response, multiple evaluation scores are included: one per annotator..
Both the platform and the experiment results file format (described in more detail in Appendix A) have been designed in a model-agnostic and metric-agnostic way, in order to support analysis of diverse evaluation experiments.
Once provided with the input file, InspectorRAGet validates it for errors (e.g., missing evaluation scores) and augments it with additional statistics (such as inter-annotator agreement for human evaluations). The augmented experiment results are then used to power the platform’s frontend; a visual analytics application that enables developers and stakeholders to interactively analyze and gain insights on different aspects of the experiment, as described in the next section.
Finally, while InspectorRAGet focuses on the analysis of the experiment results, to help users employ the platform as part of the broader RAG evaluation life cycle, we have also implemented a sample experiment runner in a Python notebook, showing how to run experiments using HuggingFace assets Wolf et al. (2020) (e.g. datasets, models, and metric evaluators) and output the results in the format expected by InspectorRAGet (see Appendix B for details).
@@ 3 InspectorRAGet Platform
InspectorRAGet’s frontend offers a series of views (presented as separate tabs), each tailored to a different aspect of the evaluation process for analysis (see Appendix C for implementation details). Excerpts of these views are shown in Figure 2. We describe each view along with the analyses it enables. In Section 4, we then demonstrate how these views can be used to extract actionable insights for two different use cases.
@@ 3.1 Predictions Table
When analyzing an experiment, it helps to first see a few examples of instances. To this end, InspectorRAGet’s predictions view shows a table of all questions with the corresponding model responses. This view not only gets developers acquainted with the experiment but also helps them spot patterns in the model responses (e.g., length, repetitions of text, etc.) that should be investigated further.
@@ 3.2 Performance Overview
After getting acquainted with the experiment, developers and stakeholders can get an overview of the experiment results through the overall performance view. This view shows the aggregate score and ranking of each model for each evaluation metric. This information is rendered both in tabular form, as well as through a Radar chart with separate tables/visualizations for human and algorithmic metrics (Figures 2b and 2c).
While designing this view, special attention was given to quantifying the uncertainty present in human evaluations to avoid misinterpretation of the results. Aggregate evaluation scores for human metrics are shown along with (a) their standard deviation and (b) a visualization of inter-annotator agreement (displayed through sparkline charts).
@@ 3.3 Model Behavior
After obtaining an overview of aggregate model performance, one can drill down and inspect model responses for individual instances through the model behavior view. Users start the analysis by filtering instances based on criteria, such as dataset domain, whether a question is answerable, etc. The view then shows a histogram of model scores for all instances satisfying the filter (Figure 2d), as well as a sortable table of the actual instances (Figure 2e). Upon selecting an instance, users see a detailed view of the instance, including the conversation/question, context, the model responses and their respective evaluation scores (Figure 2f). Instance-level analysis is crucial for conducting error analysis and understanding the root causes of issues observed in other views, as well as identifying “I know it when I see it" types of issues. The view also provides users with the functionality to easily copy, flag or comment on instances.
@@ 3.4 Model Comparator
In addition to inspecting individual model responses, one can also analyze entire models. This is enabled by the model comparator view, which for a chosen pair of models and a metric, shows whether the scores of the two models for the selected metric are derived from the same distribution (Figure 2g). This is shown both through a scatter plot, depicting the scores of the two models for the metric, as well as through the result of a statistical significance test, computed using Fisher’s randomization method (Smucker et al., 2007). Continuing the support for instance-level analysis, the view also allows one to drill down and inspect the instances where two models received very similar/dissimilar evaluation scores. This view can be useful for spotting similarities and differences between models.
@@ 3.5 Metric Behavior
Similarly to comparing models, one can also compare metrics. This is accomplished through the metric behavior view, which shows the Spearman correlation scores for each pair of metrics. This allows developers to gain insights on metrics, such as identifying whether an automatic metric correlates well with human judgements or whether supposedly orthogonal metrics correlate with each other (see fluency and answer relevance in Figure 2h), hinting at issues with metric definitions.
@@ 3.6 Annotator Behavior
Finally, when human evaluation is involved, it is imperative to also analyze its quality. This is enabled by the annotator behavior view, which contains two types of visualizations related to annotation quality: (a) Model-level visualizations that show the annotator agreement when evaluating each model (Figure 2i). These can help provide insights on challenges that certain models pose to human annotators (e.g., if humans had trouble reaching agreement when evaluating certain models). (b) Annotator-level visualizations depicting individual annotators’ performance. These include a visualization of inter-annotator agreement (Figure 2j), computed using Cohen’s kappa Cohen (1960)), annotator contribution (i.e., how often an annotator agreed with the majority) (Figure 2k), and annotation time (not shown for space reasons). Insights drawn from this view can drive additional analyses (e.g., to understand why annotators had trouble annotating the responses of certain models) or changes to the experiment setting (e.g., give feedback to individual annotators or improve the annotation guidelines).
@@
RAG Model Evaluation | CLAPnq | Llama-13B Touvron et al. (2023), GPT-3.5 Turbo Brown et al. (2020), Mistral Jiang et al. (2023) |
@@
LLM-as-a-Judge Evaluation | MT-Bench | Alpaca-13B Taori et al. (2023), Claude-v1 cla (2023), GPT-3.5 Brown et al. (2020), GPT-4 OpenAI et al. (2024), Vicuna-13B Chiang et al. (2023) |
@@
4 Finding Insights
To showcase the value of InspectorRAGet, we next describe how it can be used to get insights on two use cases: (1) RAG Model Evaluation: We collected human judgements of model responses for a RAG dataset. This allows us to show the full scope of our platform for comparing human and algorithmic metrics as well as multiple annotators. (2) LLM-as-a-Judge Evaluation: We use annotations comparing human and LLM-as-a-judge on model output for multi-turn question answering.
The experiment setting for each use case (incl. dataset, models, and evaluation metrics) is summarized in Table 1. For each use case we describe the discovered insights along with the Source views used to identify them and propose possible Actions for improving the RAG life cycle. We provide platform screenshots of the findings in Appendix D.
4.1 RAG Model Evaluation Use Case
@@
CLAPnq is built on the portion of the Natural Questions dataset Kwiatkowski et al. (2019) that only has a long answer (gold passage) without an extractive short answer. The responses in CLAPnq are grounded on the gold passage and must be concise and complete. We ensure that every question is evaluated by the same algorithmic metrics and the same number of human evaluators (3 per question). The human evaluation annotation task was completed on Appen333https://www.appen.com/, a crowdsourcing platform for collecting high-quality annotations. Each task has a question, grounding passage, and multiple randomly shown model responses for the annotator to provide their evaluations. We asked annotators to evaluate the answers on three metrics: fluency, answer relevance, and faithfulness as commonly used in the literature Es et al. (2023); Chiang and Lee (2023). In addition, we also asked them to perform a head-to-head comparison of all models for win-rate. These annotations allow us to use the platform to compare and draw insights from the models, data, metrics, and annotators. We next describe key insights we found using the platform.
@@
Dataset Inconsistencies: During the human evaluation the reference responses (labeled Reference) was rated highest by the annotators which shows that the dataset is of good quality. However, the annotators disagree most on these responses. Filtering to see examples of disagreement, shows that the most disagreement is 25 cases of Answer relevance. Opening up an indeterminate example where there was no agreement, shows a case of a tricky question+answer which explains the disagreement.
Source: Performance Overview, Annotator Behavior, Model Behavior. Action: Propose revisiting the experiment design via data analysis and improvement.
Data Types: It is also interesting to explore the answerable and unanswerable splits separately. Examples of insights include finding examples where an unanswerable question was answered by the models and is faithful, which may indicate that the question is actually answerable. One may also search for unanswerable examples that are not faithful to show that the model is coming up with an answer from its own knowledge or is hallucinating. Source: Performance Overview, Model Behavior. Action: Performance can differ due to question type, domain and other characteristics. Provide model feedback to stakeholders per dimension.
@@
4.2 LLM-as-a-Judge Evaluation Use Case
There has been a significant uptick in the popularity of LLMs as judges to evaluate the quality of LLM responses in RAG settings, complemented by active research in the efficacy of these judges Chiang and Lee (2023); Shen et al. (2023). Our platform also seamlessly supports analyzing the performance of LLM-as-a-judge approaches, and we illustrate this use case using MT-Bench. MT-Bench Zheng et al. (2023) is a multi-turn question answering dataset of 80 high quality multi-turn questions. The paper introducing this dataset explores using LLM-as-a-judge as a means of evaluating benchmarks by comparing it to human evaluation by 58 expert annotators. To this end, they release human and algorithmic GPT-4 judgements on the MT-bench dataset.
This evaluation is not centered around comparing model performance, but rather judge performance.
They discuss several limitations of LLM-as-a-judge: positional bias, verbosity bias, self-enhancement bias (judge favors its own model’s answers), and limited reasoning ability (low reasoning and math capability). They show that GPT-4 judge matches human evaluation at over 80%. We expand on their insights and introduce additional ones.

### v2

@@ Abstract
Large Language Models (LLM) have become a popular approach for implementing Retrieval-Augmented Generation (RAG) systems, and a significant amount of effort has been spent on building good models and metrics. In spite of increased recognition of the need for rigorous evaluation of RAG systems, few tools exist that go beyond the creation of model output and automatic calculation. We present InspectorRAGet, an introspection platform for performing a comprehensive analysis of the quality of RAG system output. InspectorRAGet allows the user to analyze aggregate and instance-level performance of RAG systems, using both human and algorithmic metrics as well as annotator quality. InspectorRAGet is suitable for multiple use cases and is available publicly to the community1. A live instance of the platform is available at https://ibm.biz/InspectorRAGet.
@@ 1 Introduction
There has been increased recognition of the importance of RAG evaluation Longpre et al. (2024). This consists of 1) designing, 2) running, and 3) analyzing experiments (see Section 2). The research community has mainly limited analyzing to aggregate metrics via benchmark datasets Liu et al. (2023); Chen et al. (2023), evaluation metrics Es et al. (2023) as well as evaluation frameworks (such as RAGAs Es et al. (2023) and ARES Saad-Falcon et al. (2023)).
For any input (e.g., a question), a RAG system runs a retriever and passes the input and retrieved passages to a generator. This RAG output can then be evaluated on a variety of metrics. The output is affected by variations in retriever and generator models as well as different model configurations. Improving system performance requires actionable analysis of these variations, specific to RAG. We present a platform for performing a comprehensive analysis of the quality of RAG output:
Holistic analysis: End-to-end analysis of RAG system performance involves aggregate evaluations along several dimensions and configurations, enabling continuous benchmarking of models and datasets Gehrmann et al. (2022). A comprehensive set of RAG evaluation metrics comprises algorithmic scores, LLM judges Es et al. (2023) and human judgements, and our platform enables comparison and correlation analysis to create a full picture of performance.
Quality analysis of all aspects of a RAG experiment: While most evaluation works focus on the quality of model outputs, it is important to also analyze the quality of all other aspects of a RAG experiment. These include (a) the annotator behavior (e.g. inter-annotator agreement and underperforming annotators) (b) the effectiveness of evaluation metrics at modeling the desired functionality (e.g., faithfulness to context(s)), and (c) the RAG-specific properties of the employed dataset (e.g., relevant passages and question attributes). This analysis can be used to improve the experiment as well as providing context for the interpretation of the results.
Error analysis through instance-level inspection: Actionable error analysis for RAG requires going beyond the typical aggregate-level statistics into inspection of individual instances. This is essential for developers to identify the source of undesirable output, correct erroneous reference answers, clarify ambiguous instances, and form hypotheses for model improvement.
Currently, these analyses can only be done piecemeal, requiring the manual examination of evaluation output using various ad-hoc data processing scripts and spreadsheet tools which are not sustainable, lack re-usability, and are difficult to interpret. We provide the first platform with end-to-end capabilities for thorough RAG analysis, an undersupported aspect of RAG system development intended for researchers, model developers, and stakeholders. Our contributions are as follows:
@@
We present InspectorRAGet, a rich interactive platform for performing a comprehensive analysis of RAG system output quality.
@@
We evaluate InspectorRAGet on its ability to yield concrete and actionable insights on two use cases, on both new and existing datasets.
@@
We open source the platform to the community111https://github.com/IBM/InspectorRAGet and host it on HuggingFace222https://ibm.biz/InspectorRAGet.
@@ 2 The RAG Evaluation Life Cycle
We briefly describe the life cycle of evaluating RAG systems, and situating InspectorRAGet within it, as shown in Figure 1. Evaluating RAG systems involves three main steps:
(1) Design the evaluation experiment, which consists of Models, Datasets, Metrics and Metric Evaluators, as defined below.
(2) Run the evaluation experiment and compute the evaluation scores. This involves (a) generating the model responses/output and (b) passing the responses to the metric evaluators (algorithms or humans) to produce instance-level scores.
(3) Analyze the experiment results to gain actionable insights about the models, datasets, metrics, annotation quality, etc.
@@ 2.1 InspectorRAGet for RAG Evaluation
InspectorRAGet focuses on the third step of the RAG evaluation life cycle. To analyze an experiment using the platform, users upload the standardized experiment results JSON file which contains the following information:
@@
• Metric metadata: Metadata about the metrics on which model responses were evaluated. These include metric name, type (i.e., algorithmic, LLM-as-a-judge and human, such as crowd workers), and scale (yes/no, Likert scale, numeric, etc.).
• Model output and Evaluation scores: The model responses and evaluation scores for the models along each metric on every data instance 333In case of multiple annotators per metric, such as when multiple humans annotate a single response, multiple evaluation scores are included: one per annotator..
The platform and the experiment results file format have been designed in a model-agnostic and metric-agnostic way to support analysis of diverse evaluation experiments (see Appendix A).
Once provided with the input file, InspectorRAGet validates it for errors (e.g., missing evaluation scores) and augments it with additional statistics (e.g. inter-annotator agreement for human evaluations). The augmented results power the platform’s frontend: a visual analytics application that enables users to interactively analyze and gain insights on different aspects of the experiment.
Finally, to help users employ the platform as part of the broader RAG evaluation life cycle we also provide sample experiment runners, showing how to run experiments using popular evaluation frameworks and output the results in the format expected by InspectorRAGet for analysis of experiment results. Our experiment runners showcase integrations with the Language Model Evaluation Harness Gao et al. (2023) and RAGAs Es et al. (2023) evaluation frameworks, as well as HuggingFace Wolf et al. (2020) (see Appendix B).
@@ 3 InspectorRAGet Platform
InspectorRAGet is a React web application built with NextJS 13 framework444https://react.dev, https://nextjs.org. We use the Carbon Design System555https://carbondesignsystem.com for the user interface. The experiment results are provided as a json input file that is loaded on the platform. This enables our platform to be lightweight; it can easily be run on virtual machines or even personal laptops/desktops. To enable privacy, InspectorRAGet is a stateless application and does not retain any uploaded datasets. InspectorRAGet’s frontend offers a series of views (presented as separate tabs), each tailored to a different aspect of the analysis process. Excerpts of these views are shown in Figure 2. We describe each view along with hypotheses that can be formed, analyses it enables, and insights which may require further investigation.
3.1 Dataset Characteristics
The Dataset Characteristics view (Figure 2a) provides details regarding characteristics of the dataset such as answerability (e.g. answerable, unanswerable, partial), question type (e.g. factoid, comparative, explanation), and question length. These details can provide a quick snapshot of dataset trends, outliers, and potential biases.
@@ 3.2 Predictions Table
When analyzing an experiment, it helps to first see a few examples of instances. InspectorRAGet’s predictions view shows a table of all questions with the respective model responses (Figure 2b). This view not only gets developers acquainted with the experiment but also helps them spot patterns in model responses (e.g., length, repetitions of text).
@@ 3.3 Performance Overview
After getting acquainted with the experiment, developers and stakeholders can get an overview of the experiment results through the overall performance view. This view shows the aggregate score and ranking of each model for each evaluation metric. This information is rendered both in tabular form, as well as through a Radar chart with separate tables/visualizations for human and algorithmic metrics (Figures 2c and 2d).
While designing this view, special attention was given to quantifying the uncertainty present in human evaluations to avoid misinterpretation of the results. Aggregate evaluation scores for human metrics are shown along with (a) their standard deviation and (b) a visualization of inter-annotator agreement (through sparkline charts). Despite its resemblance to leaderboards Zheng et al. (2023); Hendrycks et al. (2021), the goal of this view is not only to declare winners but to also make initial observations that need to be further explored.
@@ 3.4 Model Behavior
After obtaining an overview of aggregate model performance, one can drill down and inspect model responses for individual instances through the model behavior view. Users start the analysis by filtering instances based on criteria, such as dataset domain, whether a question is answerable, etc. The view then shows a histogram of model scores for all instances satisfying the filter (Figure 2e), as well as a sortable table of the actual instances (Figure 2f). Upon selecting an instance, users see a detailed view of the instance, including the conversation/question, context(s), the model responses and their respective evaluation scores (Figure 2g). Instance-level analysis is crucial for conducting error analysis and understanding the root causes of issues observed in other views, as well as identifying “I know it when I see it" types of issues. The view also provides users with the functionality to easily copy, flag or comment on instances.
@@ 3.5 Model Comparator
In addition to inspecting individual model responses, one can also analyze entire models. This is enabled by the model comparator view, which for a chosen pair of models and a metric, shows whether the scores of the two models for the selected metric are derived from the same distribution (Figure 2h). This is shown both through a scatter plot, depicting the scores of the two models for the metric, as well as through the result of a statistical significance test, computed using Fisher’s randomization method (Smucker et al., 2007). Continuing the support for instance-level analysis, the view also allows one to drill down and inspect the instances where two models received very similar/dissimilar evaluation scores. This view can be useful for spotting similarities and differences between models.
@@ 3.6 Metric Behavior
Similarly to comparing models, one can also compare metrics. This is accomplished through the metric behavior view, which shows the Spearman correlation scores for each pair of metrics. This allows developers to gain insights on metrics, such as identifying whether an automatic metric correlates well with human judgements or whether supposedly orthogonal metrics correlate with each other (see fluency and answer relevance in Figure 2i), hinting at issues with metric definitions.
@@ 3.7 Annotator Behavior
Finally, when human evaluation is performed, it is imperative to also analyze its quality. This is enabled by the annotator behavior view, which contains two types of visualizations related to annotation quality: (a) Model-level visualizations that show the annotator agreement when evaluating each model (Figure 2j). These can help provide insights on challenges that certain models pose to human annotators (e.g., if humans had trouble reaching agreement when evaluating certain models). (b) Annotator-level visualizations depicting individual annotators’ performance. These include a visualization of inter-annotator agreement (Figure 2k), computed using Cohen’s kappa Cohen (1960)), annotator contribution (i.e., how often an annotator agreed with the majority) (Figure 2l), and annotation time (not shown for space reasons). Insights drawn from this view can drive additional analyses (e.g., to understand why annotators had trouble annotating the responses of certain models) or changes to the experiment setting (e.g., give feedback to individual annotators or improve the annotation guidelines).
@@
| RAG Model Performance | CLAPnq | Llama-13B Touvron et al. (2023), GPT-3.5 Turbo Brown et al. (2020), Mistral Jiang et al. (2023) |
@@
| LLM-as-a-Judge Performance | MT-Bench | Alpaca-13B Taori et al. (2023), Claude-v1 Anthropic (2023), GPT-3.5 Brown et al. (2020), GPT-4 OpenAI et al. (2024), Vicuna-13B Chiang et al. (2023) |
@@
4 Evaluating InspectorRAGet
To showcase the value of InspectorRAGet, we evaluate its ability to yield actionable insights on two use cases: (1) Analyzing RAG Model Performance: We collected human judgements of model responses for a RAG dataset. This allows us to show the full scope of our platform for comparing human and algorithmic metrics as well as multiple annotators. (2) Analyzing LLM-as-a-Judge Performance: We use annotations comparing human and LLM-as-a-judge on model output for multi-turn question answering.
The experiment setting for each use case666The use cases are at ibm.biz/InspectorRAGet (e.g. dataset, models, and evaluation metrics) is summarized in Table 1. For each use case we describe the discovered insights along with the Source views used to identify them and propose possible Actions for improving the RAG experience.
4.1 Analyzing RAG Model Performance
@@
CLAPnq is built on the portion of the Natural Questions dataset Kwiatkowski et al. (2019) that only has a long answer (gold passage) without an extractive short answer. The responses in CLAPnq are grounded on the gold passage and must be concise and complete. We ensure that every question is evaluated by the same algorithmic metrics and the same number of human evaluators (3 per question). The human evaluation annotation task was completed on Appen777https://www.appen.com/, a crowdsourcing platform for collecting high-quality annotations. Each task has a question, grounding passage, and multiple randomly shown model responses for the annotator to provide their evaluations. We asked annotators to evaluate the answers on three metrics: fluency, answer relevance, and faithfulness as commonly used in the literature Es et al. (2023); Chiang and Lee (2023). In addition, we also asked them to perform a head-to-head comparison of all models for win-rate. These annotations allow us to use the platform to compare and draw insights from the models, data, metrics, and annotators. We next describe key insights we found using the platform.
@@
Dataset Inconsistencies: During the human evaluation the reference responses (labeled Reference) were rated highest by the annotators which shows that the dataset is of good quality. However, the annotators disagree most on these responses. Filtering to see specific instances of disagreement, shows that the most disagreement is 25 cases of Answer relevance. Opening up an example where there was no agreement, shows a case of a tricky question+answer which explains the disagreement. Source: Performance Overview, Annotator Behavior, Model Behavior. Action: Investigate whether data inconsistencies should be improved or removed, and if further annotator training is needed.
Data Types: It is also interesting to explore the answerable and unanswerable splits separately. Examples of insights include finding examples where an unanswerable question was answered by the models and is faithful, which may indicate that the question is actually answerable. One may also search for unanswerable examples that are not faithful to show that the model is coming up with an answer from its own knowledge or is hallucinating. Source: Performance Overview, Model Behavior, Dataset Characteristics. Action: Performance can differ due to question type, domain and other characteristics. Provide model feedback to stakeholders per dimension.
@@
4.2 Analyzing LLM-as-a-Judge Performance
There has been a significant uptick in the popularity of LLMs as judges to evaluate the quality of LLM responses in RAG settings, complemented by active research in the efficacy of these judges Chiang and Lee (2023); Shen et al. (2023). Our platform also seamlessly supports analyzing the performance of LLM-as-a-judge approaches, and we illustrate this use case using MT-Bench Zheng et al. (2023), a dataset of 80 high quality multi-turn questions. The MT-Bench authors compare LLM-as-a-judge with humans by releasing human judgments (provided by 58 experts) and algorithmic GPT-4 judgments on the MT-Bench dataset.
This InspectorRAGet use case is not centered around comparing model performance, but rather judge performance. The MT-Bench authors discuss several limitations of LLM-as-a-judge: positional bias, verbosity bias, self-enhancement bias (judge favors its own model’s answers), and limited reasoning ability (low reasoning and math capability). They show that GPT-4 judge matches human evaluation at over 80%. We expand on their insights and introduce additional ones.